# The OATMEAL Survey. II. The 3D spin-orbit obliquity of an eccentric transiting brown dwarf in the Ruprecht 147 open cluster

### v1

@@ 4.3 Evidence against primordial misalignment
A separate mechanism that describes the realignment of planetary companions is known as “resonance locking” (M. G. Witte & G. J. Savonije, 1999; G. J. Savonije, 2008). This is when the orbital motion of the companion couples to the host star’s gravity mode (g-mode) oscillation frequency, resulting in a gradual relative alignment of the companion’s orbit and star’s spin axis. J. J. Zanazzi et al. (2024) simulate a resonance lock scenario using a sunlike star and a $1\,M_{\rm J}$ companion to explore different timescales over which the companion’s spin-orbit obliquity would realign from misaligned initial conditions (values of $\psi$ from $45^{\circ}$ to $90^{\circ}$). One important relationship J. J. Zanazzi et al. (2024) show is between the realignment timescale and scaled semi-major axis $a/R_{\star}$, where resonance lock is not possible between a 2.5 Gyr old, $1\,M_{\sun}$ star (like CWW 89A) and a $1\,M_{\rm J}$ companion at $a/R_{\star}>10$ within 10 Gyr (Figure 15 in J. J. Zanazzi et al. (2024)). This scenario is for an initial $\psi=90^{\circ}$.
In our case, the companion almost a factor of 40 more massive, which means that the brown dwarf has an increased tidal potential energy and thus is able to act on the star at larger $a/R_{\star}$ than a Jupiter-mass companion. However, even when considering this $a/R_{\star}\gtrsim 12$, the shortest plausible timescale for realignment from $\psi=90^{\circ}$ (when extrapolating from J. J. Zanazzi et al. (2024)) is roughly 9-10 Gyr for CWW 89A, significantly longer than the 2.5 Gyr age of CWW 89. We also highlight the fact that the brown dwarf did not form in its current close orbit, meaning that the couple of its orbital motion to the g-mode oscillations were likely even weaker for a portion of the system’s main sequence lifetime so far. Given this, we find it unlikely that CWW 89A was realigned with the orbital of CWW 89Ab via resonance locking.
@@ 5 Summary
  • Given the stellar rotation period and projected stellar rotational velocity (G. Nowak et al., 2017), we find CWW 89Ab has a 3D spin-orbit obliquity of $\psi=15.1^{+15.0^{\circ}}_{-10.1}$ with the posterior distribution peaking at $\psi_{\rm peak}=3.1^{\circ}$.
  • We confirm trends described in S. Giacalone et al. (2024) and J. Rusznak et al. (2024) that all transiting brown dwarfs ($M_{b}\geq 13\,M_{\rm J}$) in single- or multi-star systems with measured spin-orbit obliquities are consistent with $\lambda=0^{\circ}$ (aligned) regardless of the primary $T_{\rm eff}$.
@@
We find the trend of massive aligned companions to continue with CWW 89A, plausibly following the orbital evolutionary mechanisms outlined in C. Petrovich (2015) and J. Rusznak et al. (2024), where a high circumstellar disk mass may have enabled the growth of the brown dwarf into a mass regime where it is resistant to dynamical kicks perpendicular to the disk plane that would have misaligned the brown dwarf’s orbit were it closer to or less than the mass of Jupiter. What would then follow is a coplanar high-eccentricity migration induced by the outer companion, assuming the outer companion is coplanar with the brown dwarf. Given the long realignment timescale $\tau_{\rm CE}\approx 10$ Gyr between the brown dwarf and primary star and the long circularization timescale $\tau_{\rm circ}\gtrsim 7.5$ Gyr relative to the system’s age of 2.8$\pm$0.6 Gyr, we may conclude that CWW 89Ab formed in an eccentric, aligned orbit. Following the conclusion from T. G. Beatty et al. (2018), if we also assume CWW 89Ab formed via core accretion, then a more complete picture of CWW 89Ab’s formation is: core accretion-driven formation and growth in an aligned, eccentric orbit within the circumstellar disk, followed by a coplanar high-eccentricity migration via its outer M-dwarf companion, CWW 89B, inward to its current architectural configuration.
@@
One key missing piece of information is the true separation of CWW 89B from CWW 89A. Future observations with Keck II/NIRC2 could provide constraints on the orbit CWW 89B. Additional high-resolution imaging could also be performed on TOI-2119 and TOI-2533 to search for outer companions and the source of the high RUWE value ($\gtrsim 1.6$) for both of these stars.
The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. This research was carried out, in part, at the Jet Propulsion Laboratory and the California Institute of Technology under a contract with the National Aeronautics and Space Administration and funded through the President’s and Director’s Research & Development Fund Program. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. The research was carried out, in part, at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004) and funded through the President’s and Director’s Research & Development Fund Program. TWC is supported by an NSF MPS-Ascend Postdoctoral Fellowship under award 2316566. D.H. acknowledges support from the Alfred P. Sloan Foundation and the National Aeronautics and Space Administration (80NSSC22K0781).

### v2

@@ 4.3 Evidence against primordial misalignment
4.3.1 Resonance locking
A separate mechanism that describes the realignment of planetary companions is known as “resonance locking” (witte1999; savonije2008). This is when the orbital motion of the companion couples to the host star’s gravity mode (g-mode) oscillation frequency, resulting in a gradual relative alignment of the companion’s orbit and star’s spin axis. zanazzi2024 simulate a resonance lock scenario using a sunlike star and a $1\,M_{\rm J}$ companion to explore different timescales over which the companion’s spin-orbit obliquity would realign from misaligned initial values of $\psi$ from $45^{\circ}$ to $90^{\circ}$. One important relationship zanazzi2024 show is between the realignment timescale and scaled semi-major axis $a/R_{\star}$, where resonance lock is not possible between a 2.5 Gyr old, $1\,M_{\sun}$ star (like CWW 89A) and a $1\,M_{\rm J}$ companion at $a/R_{\star}>10$ within 10 Gyr (Figure 15 in zanazzi2024). This scenario is for an initial $\psi=90^{\circ}$.
In our case, the companion is almost a factor of 40 more massive, which means that the brown dwarf has an increased tidal potential energy and thus is able to act on the star at larger $a/R_{\star}$ than a Jupiter-mass companion. However, even when considering the $a/R_{\star}\gtrsim 12$ for CWW 89Ab, the shortest plausible timescale for realignment from $\psi=90^{\circ}$ (when extrapolating from zanazzi2024) is roughly 9-10 Gyr for CWW 89A, significantly longer than the 2.5 Gyr age of CWW 89. We also highlight the fact that the brown dwarf did not form in its current close orbit, meaning that the coupling of its orbital motion to the g-mode oscillations were likely even weaker for a portion of the system’s main sequence lifetime. Given this, we find it unlikely that CWW 89A was realigned with the orbital of CWW 89Ab via resonance locking , but we acknowledge that, like with the CHEM framework, the limits of resonance locking has not been tested to these extremes in companion mass.
4.3.2 Secular perturbations from CWW 89B
The nature of CWW 89B’s orbital inclination has profound impacts on the potential alignment of the brown dwarf. A misaligned outer companion would likely make the BD susceptible to secular perturbations like the von Zeipel-Kozai-Lidov effect (zeipel1910; kozai1962; lidov1962) where the spin-orbit alignment $\lambda$ of the BD would oscillate between misaligned and aligned configurations driven by CWW 89B (naoz2016). This would imply that the our measurement of $\lambda$ took place during a part in this cycle when the BD’s spin-orbit angle is near-zero, leading us to misinterpret the system as primordially and perpetually aligned. Whether or not these oscillations excite only the BD’s eccentricity or both the eccentricity and spin-orbit angle is unclear without knowing CWW 89B’s orbital inclination.
@@ 5 Summary
  • Given the stellar rotation period and projected stellar rotational velocity from nowak17, we find CWW 89Ab has a 3D spin-orbit obliquity of $\psi=15.1^{+15.0^{\circ}}_{-10.1}$ with the posterior distribution peaking at $\psi_{\rm peak}=3.1^{\circ}$.
  • We confirm trends described in oatmeal1 and rusznak2024 that all transiting brown dwarfs in single- or multi-star systems with measured spin-orbit obliquities are consistent with $\lambda=0^{\circ}$ (aligned) regardless of the primary $T_{\rm eff}$.
@@
[…]
Following the conclusion from cww89a, if we also assume CWW 89Ab formed via core accretion, which is likely challenging at this mass, then a more complete picture of CWW 89Ab’s formation is: core accretion-driven formation and growth in an aligned, eccentric orbit within the circumstellar disk, followed by a coplanar high-eccentricity migration via its outer M-dwarf companion, CWW 89B, inward to its current architectural configuration.
@@
One key missing piece of information is orbital inclination of CWW 89B with respect to CWW 89A. Future observations with facilities like Keck II/NIRC2 could provide constraints on the orbit CWW 89B to determine this angle. Additional high-resolution imaging could also be performed on TOI-2119 and TOI-2533 to search for outer companions and the source of the high RUWE value ($\gtrsim 1.6$) for both of these stars.
[…]
Table 4: MIST median values and 68% confidence interval for CWW 89A. Here, $\mathcal{U}$[a,b] is the uniform prior bounded between $a$ and $b$, and $\mathcal{G}[a,b]$ is a Gaussian prior of mean $a$ and width $b$.footnotetext: See Table 3 in eastman2019 for a detailed description of all parameters11footnotetext: Time of conjunction is commonly reported as the “transit time”22footnotetext: $\rm{BJD_{TDB}}$is the target’s barycentric frame and corrects for light travel time33footnotetext: Time of minimum projected separation is a more correct “transit time”44footnotetext: At the epoch that minimizes the covariance between $T_{C}$ and Period55footnotetext: Use this to predict future transit times66footnotetext: Assumes no albedo and perfect redistribution
   Parameter Description Priors Values
Stellar Parameters:
      $M_{*}$. Mass ($M_{\sun}$). - $1.018^{+0.027}_{-0.026}$
      $R_{*}$. Radius ($R_{\sun}$). - $1.067\pm 0.022$
      $L_{*}$. Luminosity ($L_{\sun}$). - $0.973^{+0.034}_{-0.037}$
      $\rho_{*}$. Density (cgs). - $1.186^{+0.024}_{-0.022}$
      $\log{g}$. Surface gravity (cgs). - $4.391\pm 0.017$
      $T_{\rm eff}$. Effective temperature (K). $\mathcal{G}$[5699,200] $5553^{+82}_{-86}$
      $\rm Fe/H$}]. Metallicity (dex). $\mathcal{G}$[0.12,0.1] $0.067\pm 0.090$
      AgeAge. Age (Gyr). $\mathcal{G}$[2.76,0.61] $2.52\pm 0.61$
      $A_{V}$. V-band extinction (mag). $\mathcal{U}$[0,0.278] $0.262^{+0.023}_{-0.045}$
      $\sigma_{SED}$. SED photometry error scaling . - $1.13^{+0.47}_{-0.29}$
      $\varpi$. Parallax (mas). $\mathcal{G}$[3.286,0.018] $3.282\pm 0.018$
      $d$. Distance (pc). - $304.6\pm 1.7$
      $V_{\rm beta}$. Gaussian dispersion (m/s). $\mathcal{G}$[4000,1000] $4080^{+1000}_{-990}$
      $V_{\rm gamma}$. Lorentzian dispersion or Differential rotation (m/s). $\mathcal{U}$[1000,1000] (fixed) $1000.00$
      $V_{\rm zeta}$. Macroturbulence dispersion (m/s). - $3050\pm 930$
      $V_{\rm xi}$. Microturbulence dispersion (m/s). - $1030^{+880}_{-660}$
Brown dwarf Parameters: b
      $P$. Period (days). - $5.2926298^{+0.0000041}_{-0.000086}$
      $R_{P}$. Radius ($R_{\rm J}$). - $0.963^{+0.037}_{-0.033}$
      $M_{P}$. Mass ($M_{\rm J}$). - $37.25^{+1.6}_{-0.98}$
      $T_{C}$. Model Time of conjunction1,2 ($\rm{TJD_{TDB}}$). - $2457341.037551^{+0.000095}_{-0.000094}$
      $T_{0}$. Obs time of min proj sep3,4,5 ($\rm{BJD_{TDB}}$). - $2457346.329698^{+0.000094}_{-0.00011}$
      $a$. Semi-major axis (AU). - $0.06048^{+0.0013}_{-0.00079}$
      $i_{0}$. Inclination (Degrees). - $88.01^{+0.14}_{-0.13}$
      $b$. Transit impact parameter . - $0.429^{+0.013}_{-0.015}$
      $e$. Eccentricity . - $0.1892^{+0.0024}_{-0.0029}$
      $e\cos{\omega_{*}}$. . - $0.1836^{+0.0023}_{-0.0026}$
      $e\sin{\omega_{*}}$. . - $0.0450\pm 0.0032$
      $\omega_{*}$. Arg of periastron (Degrees). - $13.77^{+0.93}_{-0.89}$
      $|\lambda|$. Projected Spin-orbit alignment (Degrees). - $1.4^{+2.1}_{-2.5}$
      $v\sin{i_{\star}}$. Projected rotational velocity (m/s). - $4690^{+510}_{-350}$
      $T_{\rm eq}$. Equilibrium temp6 (K). - $1268\pm 15$
      $K$. RV semi-amplitude (m/s). - $4272^{+18}_{-19}$
      $R_{P}/R_{*}$. Radius of planet in stellar radii . - $0.09283^{+0.00051}_{-0.00057}$
      $a/R_{*}$. Semi-major axis in stellar radii . - $12.21^{+0.16}_{-0.14}$
      $\delta$. $\left(R_{P}/R_{*}\right)^{2}$ . - $0.008617^{+0.00010}_{-0.00011}$
      $\delta_{\rm Kepler}$. Transit depth in Kepler (frac). - $0.01076^{+0.00011}_{-0.000100}$
      $\tau$. In/egress transit duration (days). - $0.02409^{+0.00090}_{-0.00095}$
      $T_{14}$. Total transit duration (days). - $0.15017^{+0.00052}_{-0.00050}$
      $T_{FWHM}$. FWHM transit duration (days). - $0.12609^{+0.00067}_{-0.00063}$
      $\rho_{P}$. Density (cgs). - $23.4^{+1.6}_{-1.4}$
      $logg_{P}$. Surface gravity (cgs). - $4.764^{+0.019}_{-0.016}$
      $T_{P}$. Time of Periastron ($\rm{TJD_{TDB}}$). - $2457339.830^{+0.014}_{-0.013}$
      $T_{A}$. Time of asc node ($\rm{TJD_{TDB}}$). - $2457339.9660\pm 0.0050$
      $T_{D}$. Time of desc node ($\rm{TJD_{TDB}}$). - $2457342.7668^{+0.0087}_{-0.0097}$
      $M_{P}\sin i_{0}$. Minimum mass ($M_{\rm J}$). - $37.23^{+1.6}_{-0.99}$
      $M_{P}/M_{*}$. Mass ratio . - $0.03492^{+0.00053}_{-0.00076}$
Wavelength Parameters: Kepler V
      $u_{1}$. Linear limb-darkening coeff . $0.433^{+0.035}_{-0.032}$ $0.502^{+0.052}_{-0.057}$
      $u_{2}$. Quadratic limb-darkening coeff . $0.253\pm 0.047$ $0.253^{+0.054}_{-0.050}$
Telescope Parameters: FIES KPF TRES
      $\gamma_{\rm rel}$. Relative RV Offset (m/s). $45642.2^{+9.6}_{-9.0}$ $46716^{+18}_{-230}$ $3481^{+12}_{-13}$
      $\sigma_{J}$. RV Jitter (m/s). $12^{+17}_{-13}$ $5.3^{+1.8}_{-3.5}$ $6.4^{+28}_{-6.4}$
      $\sigma_{J}^{2}$. RV Jitter Variance . $170^{+750}_{-210}$ $27^{+22}_{-25}$ $40^{+1100}_{-590}$
Transit Parameters: Kepler UT 2015-10-08 (Kepler)
      $\sigma^{2}$. Added Variance . $1.972^{+0.061}_{-0.054}\times 10^{-8}$
      $F_{0}$. Baseline flux . $0.9999990\pm 0.0000023$